yomomalikebracobama:

A water park charges $5 for entry into the park and an additional $2 for each of the big water slides. Steven spent $17 on his visit to the water park. Which equation can be used to find the number of big water slides, x, that Steven went on? (1 point) a 5 + 2x = 17 b 5x + 2 = 17 c 5x + 2x = 17 d 5 + 2 = 17x

yomomalikebracobama:

ew awordproblem

yomomalikebracobama:

who tryna just gimme the thing called answer

yomomalikebracobama:

i needhelp answers help to

yomomalikebracobama:

lol

yomomalikebracobama:

@axie

yomomalikebracobama:

mine are so easy for u guys right?

xXQuintonXx:

ok, so we know that the total is 17 and its $5 to get in so, so far we can eliminate $5 from 17 so it'd be 12 and then its $2 per ride so we can do 5+2x=17

yomomalikebracobama:

-2

yomomalikebracobama:

or 5

xXQuintonXx:

it'd be A

xXQuintonXx:

because x can be 6

Joe348:

@xxquintonxx wrote:
ok, so we know that the total is 17 and its $5 to get in so, so far we can eliminate $5 from 17 so it'd be 12 and then its $2 per ride so we can do 5+2x=17
x=6 would be your final answer subtract 6 from both sides 2x=12 divide by 2 x=6

yomomalikebracobama:

IM SO SLOW IM CRY LOLOLLOLOLLLOLOLL

xXQuintonXx:

so 6 times 2=12

Joe348:

plus 5

xXQuintonXx:

omg bruh🤣

Joe348:

WHAT IM JUST STATING FACTS MAN

xXQuintonXx:

BRUH 6x2=12+5=17 x=6

Joe348:

thats what i saids man